OpenStudy (tylermckinney16):

Stacey is a sales clerk in a bicycle shop. She's paid $7.25/hour for a 38.5 hour week, plus a commission of 7% of her sales for the week. In one week Stacey's sales were $3517. Calculate Stacey's total earnings for the week. Show work.

OpenStudy (tylermckinney16):

@welshfella @mathstudent55

OpenStudy (jdoe0001):

well... we know Stacey make 7.25 per hour, and in a week she gets 38.5 hours so... she really gets paid 7.25 * 38.5 now, her commission on the sales is 7% of whatever quantity she sold we're told her sales on the week in case, is 3517 so... she made 7.25*38.5 plus commission of 7% of 3517 well, how much is 7% of 3517? we know 3517 is the 100%, what is the 7% of that? well \(\bf \begin{array}{ccllll} quantity&\% \\\hline\\ 3517&100\\ x&7 \end{array}\implies \cfrac{3517}{x}=\cfrac{100}{7}\impliedby \begin{array}{llll} \textit{solve for "x"}\\ \textit{ to get her commission} \end{array}\)

OpenStudy (tylermckinney16):

$246.19???

OpenStudy (tylermckinney16):

@jdoe0001

OpenStudy (jdoe0001):

yeap so, Stacey made that week \(\bf (7.25\cdot 38.5)\implies 279.125+commission \\ \quad \\ 279.125+246.19\)

OpenStudy (tylermckinney16):

525.315

OpenStudy (jdoe0001):

yeap

OpenStudy (tylermckinney16):

That the answer?